# Shallow IQP circuits and graph generation

### v2

@@ IV.5 Evaluation metrics
8 BP 0.24 (-0.071) 63.09 (100) 0.29 (-0.018) 52.15 (100) 0.225 (-0.0001) 69.34 (100)
ER 0.701 (-0.062) 0.0 (0.0) 0.461 (0.019) 3.1 (0.5) 0.2862 (0.066) 31.5 (51.5)
[…]
Figure 6: Scaling of the maximum mean discrepancy (MMD) across all shallow IQP models.
@@
Lastly, we quantify the bipartite structure of the generated distributions using two complementary measures. First, bipartite accuracy, defined as the percentage of bipartite graphs generated, is determined via an exact 22-coloring check. Given that this property is binary and can be very sensitive to noise—a single misplaced edge can introduce an odd-length cycle—we complement it with an expected bipartivity measure based on the spectral bipartivity index $\beta$ introduced by Estrada et al. [estradaSpectralMeasuresBipartivity2005]
@@
We compute these metrics on $512$ samples per trained model. Simulations use PennyLane’s lightning.qubit backend [bergholmPennyLaneAutomaticDifferentiation2022], while hardware experiments are executed on IBM’s $156$-qubit Aachen QPU, with each batch completing in $5.0\pm 0.5$ seconds.
@@ V Results
For each model, we perform multiple HPO experiments across different hyperparameter ranges to ensure thorough exploration. From these, we select the best-performing model for each dataset according to the following criteria: for BP datasets, the model that generates the highest proportion of bipartite graphs; and for ER datasets, the model that minimizes the expected density error.
@@ V.1 Model validation
We validate the models using noiseless $28$-qubit simulations to assess their ability to reproduce the expected density and bipartite structure. Table 2 reports the simulated results for the best-performing model on each dataset.
Across ER categories, the models reproduce average edge densities with small deviations: from $0.019$ to $0.066$. Notably, the ER–Sparse model produced $31.5\%$ bipartite graphs; since the ground-truth contained $51.5\%$ bipartites, this corresponds to a $\approx 61.1\%$ accuracy, which is comparable to the performance of BP models.
For the BP families, bipartite accuracy depends on density. To separate learned structure from random occurrence, we use the Erdős–Rényi model as a null-hypothesis baseline: for each dataset, we generate $10^{6}$ random graphs matching the target density and measure the fraction that are bipartite. Performance close to the baseline indicates random occurrence, while performance above the baseline indicates that the model captures bipartite structure beyond random generation. For all three densities, the models exceed the null-hypothesis baseline (Table 3).
@@
Dense 25.15 63.09 +37.94
Medium 32.16 52.15 +19.99
Sparse 55.82 69.34 +13.52
Overall, the BP–Sparse model attains the best combined performance, with bipartite accuracy $69.34\%$ and an empirical density error of $10^{-4}$ (Table 2). The BP-Medium and BP-Dense models reached $52.15\%$ and $63.09\%$, respectively. The comparatively strong performance of the dense model reflects its underestimation of density, which inadvertently increases the chance of generating bipartite graphs.
These results show that, in noiseless simulation, shallow IQP circuits can learn both the expected density and bipartite structure beyond a null-hypothesis baseline.
@@ V.2 Model noise resilience
We evaluate the noise resilience of our models by comparing their performance on IBM’s Aachen QPU with ideal, noiseless simulations for the $M=8$ case. Consistent outcomes across both backends indicate robustness to hardware noise, while large deviations reveal sensitivity. Through the following sections, we interchangeably use the terms NISQ and quantum hardware to refer to the QPU.
Figure 4(a) shows that the densities measured on NISQ hardware closely match the simulated values, yielding comparable outcomes across all models. Figure 5(a) further shows that the degree distributions measured on the quantum hardware remain in close agreement with the theoretical binomial at the $28$-qubit scale. Together, these findings indicate that shallow IQP models can learn and reproduce single-bodied features on both simulators and quantum hardware.
Bipartite accuracy shows substantially larger degradation on quantum hardware (Fig. 4(b)). The simulated models generated roughly twice as many bipartite graphs as NISQ executions. In contrast, the expected spectral bipartivity (Fig. 4(c)) remains similar across both backends, indicating that relaxations of binary features resist noise more effectively. The MMD comparison (Fig. 4(d)) follows a similar trend: simulations consistently outperform NISQ executions, though the absolute MMD differences remain small, below $0.05$.
Together, these observations show that the inherent noise of NISQ devices degrades the performance in proportion to each feature’s bodyness and noise sensitivity: local, single-body statistics stay robust; relaxations of global properties degrade moderately; and binary global features deteriorate the most. Crucially, at the 88-node scale, noise reduces precision but preserves the learned structure.
@@ V.3 Model scaling
Building on the validation and noise-resilience experiments, we study how shallow IQP models scale towards the limits of current quantum hardware. While the inherent imperfections of NISQ hardware limit the conclusions regarding their expressive power and potential, these results provide a reference point for their performance on current quantum devices.
We use the MMD as a global accuracy metric to track how the learned distributions diverge from the targets as the system grows (Fig. 6). Within the ER family, the ER–Medium model achieves the best performance across all sizes, showing the lowest MMD values. ER–Sparse and ER–Dense follow, maintaining MMD levels similar to BP–Sparse and BP–Medium. In contrast, the BP–Dense model performs worst at nearly all scales and degrades sharply at the $153$-qubit level.
For the ER models, we analyze local connectivity through the degree distribution and its TVD from the theoretical binomial (Fig. 5). The $28$- and $45$-qubit models reproduce the target distribution accurately, reaching TVDs below $0.055$. Larger systems yield less accurate results. At $91$ qubits, the sparse model underestimates the mean degree, generating overly sparse samples, while the medium and dense models align more closely with the target, giving an average TVD of $0.2038$. At $153$ qubits, the medium and dense models maintain similar performance, whereas the sparse model improves, reducing the average TVD to $0.1012$.
For the BP models, we examine structural properties through their generation accuracy and spectral bipartivity. Accuracy declines as connectivity and qubit count increase, dropping at higher densities and larger scales (Fig. 7), which reflects the growing complexity of this feature. At the $28$- and $45$-qubit scales, the models consistently exceed the baseline accuracy across all densities, whereas at $91$ and $153$ qubits, they generally fall below it. Spectral bipartivity follows the same trend for the smaller systems but performs slightly better overall, with the $91$-qubit models surpassing the baseline more consistently.
These results show that shallow IQP circuits can accurately reproduce low-bodied features even at high qubit counts but struggle with properties that rely on higher-order correlations.
@@ VI Conclusion
In this work, we studied shallow IQP circuits through the task of graph generation with experiments ranging from $28$-qubit simulations to $153$-qubit quantum hardware runs. Our results show that these models can effectively capture low-bodied features at scale, achieving an average TVD of $0.101$ on the generated degree distributions at $153$ qubits. In contrast, higher-bodied features proved to be more complex; however, up to the $45$-qubit scale, the models still surpassed the baseline for bipartite accuracy—the most challenging property we examined.
Furthermore, we employed one of the simplest shallow ansätze conjectured to be classically intractable and applied no error mitigation or post-processing techniques. Such methods would likely enhance performance, as our simulations indicate that reducing noise improves performance, particularly for the most noise-sensitive features. Therefore, these results establish a raw performance baseline for generative graph modeling on current quantum hardware.
Our findings indicate that shallow IQP circuits are most effective for modeling distributions dominated by low-bodied features, in particular at higher qubit counts. For global features, smaller models, despite their limitations, may be better suited as hardware noise will be less noticeable.
IQP-based generative modeling remains a young approach with significant potential. Its classical trainability enables building models that were previously not possible. However, it introduces several challenges: optimizing requires expressing the cost function in terms of expectation values; identifying suitable alternative loss functions remains an open question; and it is not known whether they suffer from barren plateaus. Future work should address these challenges. On the practical side, the impact of error mitigation and post-processing techniques should be explored, as better performance is expected. Furthermore, alternative training schemes, such as adversarial learning, could offer additional advantages and provide deeper insights into how IQP circuits function as generative models.

### v3

@@ IV.5 Evaluation metrics
8 BP 0.312 (-0.001) 49.21 (100) 0.230 (0.058) 63.28 (100) 0.228 (-0.003) 77.19 (100)
ER 0.759 (0.002) 0.0 (0.0) 0.437 (0.004) 3.9 (0.5) 0.221 (0.0005) 56.64 (51.5)
@@
Lastly, we quantify the bipartite structure of the generated distributions using two complementary measures. First, bipartite accuracy, defined as the percentage of bipartite graphs generated, is determined via an exact $2$-coloring check. Given that this property is binary and can be very sensitive to noise—a single misplaced edge can introduce an odd-length cycle—we complement it with an expected bipartivity measure based on the spectral bipartivity index $\beta$ introduced by Estrada et al. [estradaSpectralMeasuresBipartivity2005]
@@
We compute these metrics on $512$ samples per trained model. Simulations use PennyLane’s lightning.qubit backend [bergholmPennyLaneAutomaticDifferentiation2022], while hardware demonstrations are executed on IBM’s $156$-qubit Aachen QPU, with each batch completing in $5.0\pm 0.5$ seconds.
@@ V Results
For each dataset, we conduct a single HPO run and save the top $3$ best fitting models in terms of the cross-validation loss. Next, we sample to obtain the generation metrics. From these, we select the best-performing model for each dataset according to the following criteria: for BP datasets, the model that generates the highest proportion of bipartite graphs; and for ER datasets, the model that minimizes the total variation distance of the degree distribution.
@@ V.1 Model validation
In this section, we validate our approach by simulating the sampling of the $28$-qubit models in a noiseless regime. We evaluate performance based on the expected density—which corresponds to single-bodied correlations—and bipartite structure, which depends on all odd-order correlations. Together, these metrics probe the two ends of the expressivity spectrum. Successful modeling of the expected density implies the capacity to learn the lowest-order correlations. While the accuracy in capturing the bipartite structure examines whether the method is sufficiently expressive to represent complex, high-order dependencies.
All results are shown in Table 2. For ER datasets, we observe negligible deviations in expected density across all connectivity profiles, with a maximum error of $4\times 10^{-3}$. In the context of $8$-node graphs ($28$ potential edges), this corresponds to an average overestimate of just $\sim 0.1$ edges, indicating that the mean edge probability is captured with high precision. To further confirm this finding, we calculate the TVD of the degree distributions, which provides a broader assessment of the accuracy. We find that the models maintain high precision across all densities, yielding TVD values of approximately $0.0075$ (Dense), $0.0138$ (Medium), and $0.0085$ (Sparse). Collectively, these results confirm that shallow IQP models are capable of reproducing first-order correlations with high precision in a noiseless environment.
For the BP datasets, we first distinguish learned structure from random occurrence. We employ the Erdős–Rényi model as a null-hypothesis baseline: for each dataset, we generate $10^{6}$ random graphs matching the target density and measure the bipartite fraction. Performance exceeding this baseline indicates that the model can reproduce the bipartite structure beyond random chance.
@@
Dense 25.159 49.218 +24.059
Medium 32.163 63.281 +31.118
Sparse 55.822 75.195 +19.373
As shown in Table 3, models across all densities significantly outperform the null hypothesis. The BP-Sparse model attains the best combined performance, reaching a bipartite accuracy of $75.19\%$—an improvement of nearly $20$ percentage points over the baseline—while maintaining an empirical density error of $3\times 10^{-3}$ (Table 2). The BP-Medium and BP-Dense models demonstrate stronger performance on bipartite generation, achieving accuracies of $63.281\%$ and $49.218\%$ respectively, but fall shorter on the empirical density error.
Collectively, these results show that in a noiseless $28$-qubit regime, shallow IQP circuits can effectively capture both local and global correlations. The models reproduce the expected density with high precision and bipartite structure beyond the null-hypothesis baseline. For broader context, we provide a comparison with classical generative models in Appendix B.
@@ V.2 Model noise resilience
[…]
Before scaling to regimes beyond classical simulation, we assess the noise resilience of our models by benchmarking $28$-qubit performance on IBM’s Aachen QPU against noiseless simulations. Consistency between backends would signal robustness to hardware noise, whereas significant deviations highlight sensitivity. The quantum device characteristics and parameters at the time of data acquisition are outlined in Appendix C.
Figure 4(a) illustrates a near-perfect alignment between NISQ-measured densities and simulated values, confirming that single-bodied statistics are minimally affected by hardware noise at this scale. To further assess the impact of noise on such features, we compare the TVD of the degree distributions. While noiseless simulations yielded an average TVD of $\sim 0.01$, the quantum hardware achieves an average of $\sim 0.0307$ (see Fig. 5). This increase is expected due to hardware noise, yet the value remains low enough to confirm that local structures are robustly reproduced.
Remarkably, global features also exhibit notable resilience. Despite noise-induced errors, the bipartite accuracy on NISQ hardware (Fig. 4(b)) closely tracks the noiseless ideal, with data points clustering tightly along the diagonal. This suggests that the learned bipartite structure is sufficiently robust for the noise profile of the quantum hardware at the $28$-qubit scale. Similarly, the expected spectral bipartivity (Fig. 4(c)) remains in close agreement across backends. Finally, the MMD comparison (Fig. 4(d)) corroborates this high precision; deviations between the results observed on simulations and quantum hardware are constrained to the order of $10^{-3}$.
Together, these observations suggest that while hardware noise impacts performance, it is not decisive at this scale. Local, single-body statistics exhibit the highest robustness, whereas global properties are comparatively more susceptible. However, contrary to the expectation that fragile global features would rapidly decay, we find that even strict bipartiteness is preserved with remarkable consistency at this scale.
@@ V.3 Model scaling
Figure 6: Scaling of the maximum mean discrepancy (MMD) across all models. MMD values are computed between samples produced on NISQ hardware and the target dataset.
Building on the validation and noise-resilience analysis, we now investigate the scaling behavior of shallow IQP models as they approach the limits of available quantum hardware. While the inherent imperfections of the NISQ regime constrain definitive conclusions regarding expressivity, these results establish an empirical baseline for generative performance on state-of-the-art devices.
We analyze the local connectivity of the ER models via the degree distribution and its TVD from the theoretical binomial (Fig. 5). The $28$- and $45$-qubit configurations reproduce the target statistics with high fidelity, yielding average TVDs of $0.0307$ and $0.0478$, respectively. However, deviation increases with system size: average TVDs rise to $0.0898$ at $91$ qubits and $0.1339$ at $153$ qubits, representing a $\sim 3\times$ increase relative to the smallest instances. Given that the complexity of reproducing the degree distribution is independent of system size, we attribute the degradation primarily to noise. Despite this quantitative decline, qualitative performance remains consistent; as the histograms demonstrate, the generated distributions retain the characteristic binomial shape across all densities.
We employ the MMD as a global accuracy metric to track the divergence between learned and target distributions as system size increases (Fig. 6). The results reveal a clear hierarchy of learnability dictated by structural complexity. The Erdős-Rényi families (orange) prove to be the easiest to learn, maintaining consistently low MMD values (below $10^{-2}$) across all scales. In contrast, the bipartite distributions (blue) exhibit higher divergence, particularly in the dense regime.
Notably, the MMD does not exhibit a monotonic increase with system size for BP models, as we would expect from the increased complexity at larger scales. For instance, the BP-Sparse configuration shows improving MMD values at larger scales, even as global features like strict bipartiteness degrade. This counterintuitive behavior underscores a critical limitation of using the MMD as a standalone metric in this context: the statistical correlations that dominate the loss landscape do not necessarily align with the defining structural features of the target distribution. Consequently, while a low MMD is a necessary condition for convergence, it is insufficient for validation; a rigorous assessment requires inspecting specific structural features of the generated data to fully characterize the model’s fidelity.
For the BP models, we evaluate the learning of high-order correlations via the bipartite generation accuracy and spectral bipartivity (Fig. 7). Both metrics exhibit a similar trend: performance degrades significantly as system size and density increase, reflecting the difficulty of maintaining global constraints at larger scales where noise is more pronounced. Strict bipartite accuracy drops to zero for the $91$- and $153$-qubit models shortly after the baseline, whereas smaller models consistently maintain higher performance. Notably, spectral bipartivity follows a similar decay but maintains a score exceeding the random baseline even when strict bipartite accuracy is zero (e.g., $\rho\approx 0.2-0.4$ at $91$ qubits). This indicates that partial bipartite features persist even if they are insufficient to satisfy the exact global constraint. Furthermore, the $91$- and $153$-qubit models fail to generate graphs beyond a specific connectivity threshold. This behavior suggests an optimization trade-off at large scales, unable to simultaneously reproduce the target density and enforce bipartiteness; the models appear to prioritize lower-density configurations where the constraint is less restrictive, effectively failing to populate the denser regions of the target distribution.
Collectively, these results demonstrate that our approach maintains high fidelity for local features, such as degree distributions, even at the $153$-qubit scale. While global properties dependent on high-order correlations degrade with system size, the analysis of spectral bipartivity reveals that structural learning persists beyond the limits of strict bipartite accuracy. Ultimately, the models exhibit a clear hierarchy of learnability: unconstrained and local structures remain robust, whereas the combination of global features with strict constraints represents the current frontier for our approach on NISQ hardware.
@@ VI Conclusion
In this work, we have introduced a method for quantum generative modeling on NISQ devices that combines the theoretical advantages of IQP circuits with a practical implementation strategy. By exploiting the classical trainability of the IQP ansatz, our approach bypasses the sampling overhead and noise constraints during training, enabling the optimization of large-scale models. Furthermore, we minimize the resources required for implementation while retaining classical hardness through an ansatz that maps to a 2D Ising model with a magnetic field.
We validated this approach through a formalized graph generation setup, which allowed us to probe the relationship between feature locality and model accuracy on simulations and real quantum hardware, ranging from $28$ to $153$ qubits. The proposed method demonstrated exceptional robustness in capturing local features, accurately reproducing the degree distribution of target graphs up to the full $153$-qubit scale without error mitigation. In contrast, global structural constraints such as bipartiteness were successfully learned and reproduced up to $45$ qubits—outperforming random baselines—but degraded at larger scales due to hardware noise.
These results position shallow IQP circuits as a viable, scalable solution for NISQ generative modeling. The ability to learn and generate distributions defined by local and intermediate-range correlations at over $100$ qubits suggests practical significance in domains where such features dominate, as our method is not limited to graph generation.
DATA AVAILABILITY
The code and data supporting the findings of this study are openly available in Zenodo [ballo_gimbernat_2026_zenodo].
@@ Acknowledgments
Appendix A Derivation of the maximum mean discrepancy as a quantum observable
This appendix details the derivation of the squared Maximum Mean Discrepancy (MMD) as a mixture of Pauli-Z expectation values. This formulation, first introduced in Rudolph et al. 2024 [rudolphTrainabilityBarriersOpportunities2024], allows the classical training of Instantaneous Quantum Polynomial (IQP) circuits as implicit generative models.
The squared MMD is an integral probability metric that quantifies the distance between a target distribution $p$ and a model distribution $q_{\theta}$. It is defined via a kernel $k(\bm{x},\bm{y})$ as,
$\begin{split}\text{MMD}^{2}(p,q_{\theta})&=\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim p}[k(\bm{x},\bm{y})]\\ &\quad-2\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})]\\ &\quad+\mathbb{E}_{\bm{x}\sim q_{\theta},\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})].\end{split}$ (18)
Crucial to this derivation, we employ the Gaussian kernel,
[…]
Our goal is to rewrite the Euclidean distance $\bm{x}-\bm{y}||^{2}$ with quantum operators. First, since $\bm{x}$ and $\bm{y}$ are binary strings, the expression becomes the Hamming distance,
$\bm{x}-\bm{y}||^{2}=\sum_{i=1}^{n}(x_{i}-y_{i})^{2}.$ (20)
Next, through the properties of the exponential, the kernel becomes
$k(\bm{x},\bm{y})=\prod_{i=1}^{n}\exp-\frac{(x_{i}-y_{i})^{2}}{2\sigma^{2}},$ (21)
since $e^{a+b}=e^{a}e^{b}$. Then we can write the application of the kernel on a single bit as,
$k_{i}(x_{i},y_{i})=\exp-\frac{(x_{i}-y_{i})^{2}}{2\sigma^{2}},$ (22)
which is $1$ whenever $x_{i}=y_{i}$ and $e^{-1/2\sigma^{2}}$ otherwise. We seek an operator that satisfies those same conditions. In Rudolph et al. 2024 [rudolphTrainabilityBarriersOpportunities2024], they show that
$k_{i}(x_{i},y_{i})=\exp-\frac{(x_{i}-y_{i})^{2}}{2\sigma^{2}}\equiv(1-p_{\sigma})+p_{\sigma}(-1)^{x_{i}}(-1)^{y_{i}},$ (23)
where $p_{\sigma}$ is a scalar parameter that we can obtain by enforcing the previous conditions. For $x_{i}=y_{i}$ we trivially obtain $(1-p_{\sigma})+p_{\sigma}=1$. For $x_{i}\neq y_{i}$ we obtain $(1-p_{\sigma})-p_{\sigma}=e^{-1/2\sigma^{2}}$, which yields
$p_{\sigma}=\frac{1-e^{-\frac{1}{2\sigma^{2}}}}{2}.$ (24)
We observe that the term $(-1)^{x_{i}}(-1)^{y_{i}}$ corresponds to the eigenvalue of the Pauli-Z tensor product $Z_{i}\otimes Z_{i}$ acting on the computational basis states $|x_{i}\rangle\otimes|y_{i}\rangle$. Consequently, the single-bit kernel is equivalent to the expectation value of the quantum operator
$O_{i}=(1-p_{\sigma})\mathbb{I}\otimes\mathbb{I}+p_{\sigma}(Z_{i}\otimes Z_{i}).$ (25)
The full kernel operator is then the tensor product of these single-qubit operators across all n qubits,
$\hat{K}=\bigotimes_{i=1}^{n}\left[(1-p_{\sigma})\mathbb{I}\otimes\mathbb{I}+p_{\sigma}(Z_{i}\otimes Z_{i})\right].$ (26)
Expanding this tensor product yields a sum over all possible Pauli-Z strings. We define $Z_{\bm{a}}=\bigotimes_{i=1}^{n}Z_{i}^{a_{i}}$ for a bitstring $\bm{a}\in\{0,1\}^{n}$, where the operator $Z_{i}$ is applied if and only if $a_{i}=1$. The weight of each term in the expansion is determined by the number of $Z$ operators chosen (the Hamming weight $|\bm{a}|$). This results in the operator expansion:
$\hat{K}=\sum_{\bm{a}\in\{0,1\}^{n}}\mathcal{P}_{\sigma}(\bm{a})(Z_{\bm{a}}\otimes Z_{\bm{a}}),$ (27)
where $\mathcal{P}_{\sigma}(\bm{a})$ corresponds to the probability of sampling the string $\bm{a}$ from a product of Bernoulli distributions,
$\mathcal{P}_{\sigma}(\bm{a})=(1-p_{\sigma})^{n-|\bm{a}|}p_{\sigma}^{|\bm{a}|}.$ (28)
Returning to the classical domain, the kernel evaluation is given by the expectation of this operator with respect to the basis states $|\bm{x}\rangle$ and $|\bm{y}\rangle$, $\langle\bm{x},\bm{y}|\hat{K}|\bm{x},\bm{y}\rangle$. We note that the eigenvalue of the operator $Z_{\bm{a}}$ on a computational basis state is the product of the individual bitwise eigenvalues,
$Z_{\bm{a}}|\bm{x}\rangle=(-1)^{\bm{x}\cdot\bm{a}}|\bm{x}\rangle$ (29)
Substituting this eigenvalue into the operator expansion, the kernel function becomes,
$k(\bm{x},\bm{y})=\sum_{\bm{a}\in\{0,1\}^{n}}\mathcal{P}_{\sigma}(\bm{a})(-1)^{\bm{x}\cdot\bm{a}}(-1)^{\bm{y}\cdot\bm{a}}.$ (30)
Finally, we substitute this expansion into the definition of the MMD. By linearity of expectation, we can interchange the sum and the expectations. For the cross-term, this yields
$\mathbb{E}_{\bm{x}\sim p,\bm{y}\sim q_{\theta}}[k(\bm{x},\bm{y})]=\sum_{\bm{a}}\mathcal{P}_{\sigma}(\bm{a})\mathbb{E}_{\bm{x}\sim p}[(-1)^{\bm{x}\cdot\bm{a}}]\mathbb{E}_{\bm{y}\sim q_{\theta}}[(-1)^{\bm{y}\cdot\bm{a}}].$ (31)
Identifying the inner expectations as the moments $\langle Z_{\bm{a}}\rangle_{p}$ and $\langle Z_{\bm{a}}\rangle_{q_{\theta}}$ respectively, and applying the same logic to the self-similarity terms, the MMD simplifies to a weighted sum of squared differences between expectations,
$\text{MMD}^{2}(p,q_{\theta})=\mathbb{E}_{\bm{a}\sim\mathcal{P}_{\sigma}}\left[\left(\langle Z_{\bm{a}}\rangle_{p}-\langle Z_{\bm{a}}\rangle_{q_{\theta}}\right)^{2}\right].$ (32)
Appendix B Comparison with classical models
The classical training of IQP circuits naturally raises questions regarding whether this hybrid approach compromises potential quantum advantages or if it simply reproduces results achievable by standard classical generative models. To address this, we compare the performance of the proposed IQP models against a GraphVAE [simonovskyGraphVAEGenerationSmall2018] baseline for both bipartite and Erdős–Rényi (ER) graph generation.
B.1 GraphVAE Architecture
The GraphVAE [simonovskyGraphVAEGenerationSmall2018] is a deep generative model specifically designed to generate small graphs (between $9$ and $38$ nodes). It formulates the generation process within a Variational Autoencoder (VAE) [kingmaIntroductionVariationalAutoencoders2019] that outputs a probabilistic, fully-connected graph of a predefined maximum size in a single shot, similar to the sampling of quantum circuits.
The model consists of a graph encoder $q_{\phi}(\mathbf{z}|G)$ and a deterministic decoder $p_{\theta}(G|\mathbf{z})$. While the original architecture supports graphs with variable node counts and features $H$, we simplify the model for our fixed-size, unweighted, undirected datasets. We omit node feature vectors and, since the node count is fixed to $N$, we remove the node-existence predictions (originally encoded on the diagonal) by enforcing a zero diagonal ($\tilde{A}_{ii}=0$) in the output adjacency matrix $A\in\{0,1\}^{N\times N}$.
The encoder utilizes a feed-forward network employing Edge-Conditioned Graph Convolutions (ECC) [simonovskyGraphVAEGenerationSmall2018] to embed the input graph into a continuous latent representation. This process propagates node information through the graph layers, which is then aggregated into a fixed-size graph representation using a gated global pooling mechanism. This aggregated representation parameterizes the variational posterior $q_{\phi}(\mathbf{z}|G)$ (assumed to be a multivariate Gaussian), from which the latent vector $\mathbf{z}\in\mathbb{R}^{d_{z}}$ is sampled via the reparameterization trick.
The decoder is structured as a multi-layer perceptron (MLP) that maps the latent vector $\mathbf{z}$ directly to a probabilistic adjacency matrix $\tilde{A}\in[0,1]^{N\times N}$. The entries $\tilde{A}_{i,j}$ represent the probability of an edge existing between nodes $i$ and $j$. The model assumes that these edge probabilities are Bernoulli variables that are independent conditioned on $\mathbf{z}$. Crucially, this independence assumption does not limit the model’s ability to capture global structure (such as bipartiteness). The complex correlations required for such global features are encoded within the latent variable $\mathbf{z}$ itself; the decoder simply acts as a mapping from this globally correlated latent state to the individual edge probabilities.
B.2 Training Objective and Graph Matching
A fundamental challenge in graph generation tackled by this model is the lack of a canonical node ordering; a generated adjacency matrix $\hat{A}$ may be isomorphic to the ground truth $A$ under a node permutation $P$, yet appear distinct in a standard element-wise comparison. To evaluate the reconstruction likelihood $p_{\theta}(G|\mathbf{z})$, the model must align the generated probabilistic graph with the ground truth.
The GraphVAE optimizes an Evidence Lower Bound (ELBO) objective, comprising a reconstruction loss and a Kullback-Leibler (KL) divergence term. While the original architecture employs approximate graph matching (specifically Max-Pooling Matching) to scale to larger graphs, the small system size considered in this comparison ($N=8$) allows for tractable exact matching. We define the reconstruction loss by explicitly iterating over the set of all possible permutation matrices $\mathcal{P}_{N}$ to find the optimal alignment,
$\mathcal{L}_{\text{recon}}=\min_{P\in\mathcal{P}_{N}}\mathcal{L}_{\text{BCE}}(\hat{A},PAP^{T}),$ (33)
where $\mathcal{L}_{\text{BCE}}$ denotes the binary cross-entropy loss. This ensures that the training objective is invariant to the arbitrary node ordering of the dataset.
B.3 Results
The GraphVAE models were subjected to hyperparameter optimization (HPO) with a computational budget comparable to that of the shallow IQP circuits. We define this budget as the total number of optimization steps across all HPO trials (approximately $10^{4}$ total epochs for both model classes). The GraphVAE search space swept over the latent space dimension $d_{z}\in[16,64]$, learning rate $\eta\in[10^{-4},10^{-2}]$, and the KL annealing rate $\beta\in[10^{-3},1.0]$. Table 4 presents the performance comparison for bipartite generation accuracy. While the GraphVAE models generally achieve higher accuracy in capturing the global bipartite structure—benefiting from their high parameter count and deep inductive biases—shallow IQP circuits remain competitive, particularly given the orders-of-magnitude difference in model complexity.
Table 4: Comparison of bipartite generation accuracy between shallow IQP circuits and GraphVAE models trained on 88-node bipartite datasets.
Density Model Bip (%)
Sparse GraphVAE 86.5
IQP 75.19
Medium GraphVAE 86.1
IQP 63.3
Dense GraphVAE 80.9
IQP 49.2
Interestingly, for Erdős–Rényi (ER) datasets, the performance gap narrows significantly. As shown in Table 5, shallow IQP circuits outperform the GraphVAE baseline in reproducing the target average edge probability $\rho$ in all three cases.
Table 5: Comparison of generated average edge probability between shallow IQP circuits and GraphVAE models trained on 88-node ER datasets.
Density Model $\rho$ Target $\rho$
Sparse GraphVAE 0.2242 0.2207
IQP 0.2204
Medium GraphVAE 0.4308 0.4420
IQP 0.4419
Dense GraphVAE 0.7543 0.7618
IQP 0.7617
To contextualize these results, Table 6 compares the parameter counts across models against the performance in terms of the Total Variation Distance (TVD) of the generated degree distributions. The shallow IQP circuits achieve comparable or superior performance on these local features while utilizing approximately $3,500\times$ fewer parameters ($56$ vs. $\approx 200,000$).
Table 6: Comparison of parameter count and TVD of the generated degree distributions across models trained on 88-node datasets.
Graph Density Model Params TVD
Bipartite Sparse GraphVAE 205.2k 0.0811
IQP 56 0.0189
Medium GraphVAE 200.1k 0.0629
IQP 56 0.1428
Dense GraphVAE 199.4k 0.0789
IQP 56 0.0357
ER Sparse GraphVAE 200.1k 0.0242
IQP 56 0.0085
Medium GraphVAE 202.8k 0.0597
IQP 56 0.0137
Dense GraphVAE 206.7k 0.0382
IQP 56 0.0074
We emphasize that this comparison is not intended to demonstrate “quantum advantage” in the strict computational sense. Rather, it serves as an empirical benchmark to contextualize our approach against a classical model designed for a similar task. A key takeaway from these results is the parameter efficiency. The quantum model achieves competitive performance with orders of magnitude fewer parameters. However, we acknowledge that with larger datasets or an increased computational budget, the performance landscape could shift; however, for the tasks considered here, the shallow IQP approach offers a remarkably efficient alternative.
Appendix C Quantum device characteristics and demonstration parameters
The quantum demonstrations reported in this work were executed on the ibm_aachen quantum processing unit (QPU), a superconducting quantum computing platform accessed via the IBM Quantum cloud service. In accordance with the guidelines for cloud quantum computing demonstrations, we provide the hardware characteristics at the time of execution.
C.1 Device layout and connectivity
The ibm_aachen device features a heavy-hexagonal lattice topology. The specific physical layout is illustrated in Fig. 8.
Figure 8: Topology and connectivity graph of the ibm_aachen QPU at the time of the demonstration. Nodes represent physical qubits, with the color map indicating readout assignment error. Edges denote the RZZ error. In both cases, darker color signifies smaller errors.
C.2 Coherence and gate fidelities
Device parameters fluctuate over time; the values presented here represent the state of the QPU during our data acquisition. Table 7 summarizes the key performance metrics across the active qubits, extracted from the device calibration data provided by the backend. Specific qubit frequencies were not available in the retrieved calibration payload and are thus omitted.
Table 7: Summary of ibm_aachen device parameters during the demonstration. Values represent the range, median, and mean extracted from the device calibration dataset.
Parameter Range Median Mean
Relaxation time, $T_{1}$ (µs) 2.46499.572.46-499.57 $219.10$ $220.12$
Dephasing time, $T_{2}$ (µs) 3.35485.633.35-485.63 $182.58$ $189.25$
Single-qubit gate error $1.0\times 10^{-4}-3.18\times 10^{-2}$ $2.0\times 10^{-4}$ $7.0\times 10^{-4}$
Two-qubit (CZ) gate error $9.0\times 10^{-4}-1.91\times 10^{-1}$ $2.0\times 10^{-3}$ $1.05\times 10^{-2}$
Readout assignment error 0.00170.50000.0017-0.5000 $0.0108$ $0.0681$
Prob meas0 prep1 0.00001.00000.0000-1.0000 $0.0132$ $0.0836$
Prob meas1 prep0 0.00001.00000.0000-1.0000 $0.0045$ $0.0534$
Single-qubit gate length (ns) $32$ $32$ $32$
Readout length (ns) $2600$ $2600$ $2600$
C.3 Data analysis and mitigation
For this demonstration, the raw bitstrings obtained from the ibm_aachen QPU were used directly to evaluate the circuits. No algorithmic error mitigation, readout error mitigation, or other post-processing techniques were applied to the results. The data presented reflects the bare hardware performance under the calibrated conditions specified above.